SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (Mar.14, 2006)

Attendees: Rainer Bartoldus, Kris Chaisanguanthum, Al Eisner, Fang Fang, Selina Li, Rahmat Rahmat, Dong Su, Dan Walker

Operations

  • There is a scheduled down time for the Q2 chamber/bellow replacement from March 27 to April 4 (dates are subject to changes).
  • Rainer reported that he didn't see the hump at z~2.3m in L3 plots for the burst sample taken with the PHYSICS_BURST configuration key. This is also consistent with what Selina sees lately in these PHYSICS_BURST runs. Rainer pointed out that there are discussions in the background HN, so Selina posted new observations about the burst sample in the background HN after the meeting.
  • Al raised the question of whether the single beam runs and physics_lowlumi runs are labled correctly as a result of the hardware trigger change with SVT radiation sampling. As a reminder, we have been using the "PHYSICS_BURST" run type to take data in attempt to catch events for investigating the vacuum burst problem. Rainer mentioned that the injection samples are really SVT radiation samples with the current hardware trigger installed. Matt Weaver will probably reverting this trigger line soon.

EMT

  • The DQM shifter noted in the elog for day shift today that the EMT M clusters plot had a double peak which is different from the reference. We concluded that this is due to cosmic contamination as we are taking data with low lumi today.
  • The efficiency turn-on curves have been included in the .ps archive since last Tuesday.

GLT

  • Su Dong had a discussion with Bill, and the schedule for getting high lumi configuration will be this summer. There is room to play with logic and threshold to tighten up the trigger configuration.

L3

  • Rainer will be making a patch and then install the pending version of L3 tracking code that deals with the bug for high occupancy events. He will post to the HN when this happens.

Data Quality and Monitoring

  • Rahmat updated the L1T reference plots for fastmon in JAS today.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.